A 980kg car strikes a post at 8m/'s and comes to a stop in 0.20s. Find its initial momentum and the average force on the car while it is being stopped.


1
Expert's answer
2021-06-25T07:00:13-0400

The initial momentum is


"mv_0=980 \ kg(8\ m/s)=7840\ kg\cdot m/s"

The average force on the car is


"F_{ave}=\dfrac{mv_0}{t}=\dfrac{7840\ kg\cdot m/s}{0.20\ s}=39200\ N"